# AutoMixer: Checkpoint Artifacts as Automatic Data Mixers

### v1

@@ 1 Introduction
Therefore, the core missing piece is a direct modeling of the relationship between datasets and model parameters, because data quality cannot be reliably assessed in isolation from training (Park et al., 2023). A potential solution involves computing the influence function (Hampel, 1974; Halevy et al., 2009), which estimates the first-order “alignment” between training samples and specific skills. However, models inherently progress beyond these approximations, so data composition guided solely by step-$t$ influence calculations may become outdated as training advances.
@@ 3 The AutoMixer Framework
Concretely, this process of data regrouping is achieved by utilizing performance-based checkpoints, in which AutoMixer identifies the most effective stopping points during training (See Figure 2). The idea is that when a model achieves peak performance on task $j$ at a particular training step, the checkpoint can be employed as an effective sampler to compute influence estimation Koh et al. (2019a, b); Ting and Brochu (2018) for task $j$. Empirically, we obtain the checkpoints from simulation runs, and leverage them to identify task-optimal datasets in a two-step process (See Figure 2). The aggregated task influences can then be used to determine the weights over each group, which dictates the sampling probability of groups during language model pretraining. [^1] Suppose a data group has a weight $w_{g}=0.2$, and all weights are normalized to $1$, this means, on average, $20$% of the tokens sampled per batch will come from this group.
@@
Table 1: Checkpoint progression ratio for various tasks across different model sizes for a total of 100K steps: For instance, if the checkpoint that performs the best at a task is stored at the $5000$th step, it is then recorded on the table as 5%. Checkpoints for each model size are collected from one simulation run by training the language model on FineWeb data Lozhkov et al. (2024) with 100K steps. Here we set the tasks to be the considered benchmarks: ARC-easy, ARC-challenge Clark et al. (2018), BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag Zellers et al. (2019), OBQA Mihaylov et al. (2018), and WinoGrande Sakaguchi et al. (2021).
@@ 1.5B Parameters
Table 2: Improvements over the uniform baseline (accuracy %), each with two decimal places and averaged over 2 runs. Negative values (e.g. $-0.10$) indicate worse performance than uniform for those specific tasks. Bolded entries represent the highest improvement in each column. Overall, AutoMixer-350M yields the largest gains across most tasks, while other methods occasionally underperform vs. uniform.
@@ 8 Further Discussion
Figure 5: Mean text quality score by range: We show the sample quality score across all buckets (grouped by percentiles) of samples sorted by influence scores. The normalized mean token count (in range $[0,1]$) per sample in the same set of buckets is labeled on each point. 75M proxy model tends to select longer sentences with higher influences.

### v2

@@ 1 Introduction
[…]
@@ 3 The AutoMixer Framework
Concretely, this process of data regrouping is achieved by utilizing performance-based checkpoints, in which AutoMixer identifies the most effective stopping points during training (See Figure 2). The idea is that when a model achieves peak performance on task $j$ at a particular training step, the checkpoint can be employed as an effective sampler to compute influence estimation Koh et al. (2019a, b); Ting and Brochu (2018) for task $j$. Empirically, we obtain the checkpoints from simulation runs, and leverage them to identify task-optimal datasets in a two-step process (See Figure 2). The aggregated task influences can then be used to determine the weights over each group, which dictates the sampling probability of groups during language model pretraining. [^1] Suppose a data group has a weight $w_{g}=0.2$, and all weights are normalized to 111, this means, on average, 202020% of the tokens sampled per batch will come from this group.
@@
Table 1: Checkpoint progression ratio for various tasks across different model sizes for a total of 100K steps: For instance, if the checkpoint that performs the best at a task is stored at the 500050005000th step, it is then recorded on the table as 5%. Checkpoints for each model size are collected from one simulation run by training the language model on FineWeb data Lozhkov et al. (2024) with 100K steps. Here we set the tasks to be the considered benchmarks: ARC-easy, ARC-challenge Clark et al. (2018), BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag Zellers et al. (2019), OBQA Mihaylov et al. (2018), and WinoGrande Sakaguchi et al. (2021).
@@ 1.5B Parameters
Table 2: Improvements over the uniform baseline (accuracy %), each with two decimal places and averaged over 2 runs. Negative values (e.g. 0.10-0.10- 0.10) indicate worse performance than uniform for those specific tasks. Bolded entries represent the highest improvement in each column. Overall, AutoMixer-350M yields the largest gains across most tasks, while other methods occasionally underperform vs. uniform.
@@ 8 Further Discussion
Figure 5: Mean text quality score by range: We show the sample quality score across all buckets (grouped by percentiles) of samples sorted by influence scores. The normalized mean token count (in range [0,1][0,1][ 0 , 1 ]) per sample in the same set of buckets is labeled on each point. 75M proxy model tends to select longer sentences with higher influences.